OpenStudy (adamconner):

Please help! Medal and Fan!! -> Suppose you were to make a ring from a single strand of gold atoms. How many gold atoms would be required to make such a ring? How much would the ring weigh in grams?

OpenStudy (aaronq):

 Length of strand=Circumference=\(2*\pi*r\) \(\# \;of\;Au_{atoms}=\dfrac{length\;of\;strand}{diameter\;of \;Au_{atom}}\)

OpenStudy (aaronq):

How much would the ring weigh in grams? find moles (n): \(n=\dfrac{\#\;of\;Au_{atoms}}{N_A}\) ; \(N_A\)=avogadros number convert moles (n) to grams (m): \(n=\dfrac{m}{M}\) ; M=molar mass

OpenStudy (adamconner):

how do i get the radius of the circumference??

OpenStudy (aaronq):

idk, the didn't tell you anything about the ring? they have to give you some sort of set of starting conditions.. otherwise i'm guessing that you have to speculate and use any size (say 5 which is 49.3 mm).

OpenStudy (adamconner):

Nope. What i wrote for the question is all it says. So i will just use the 49.3 mm and ill let u know what i get.

OpenStudy (aaronq):

we'll you should ask your instructor what the size of the ring is.